# DeepDISC-photoz: Deep Learning-Based Photometric Redshift Estimation for Rubin LSST

### v1

@@ 1 Introduction
This paper is organized as follows. In §2, we describe the simulated data and pre-processing steps. In §3, we describe our model and training methodology. We present the results of our trained photo-z code in §4, and quantify our photo-z estimation performance using various metrics. In §5, we contextualize our findings and conclude.
@@ 2 Data
Figure 1: An example DC2 image used for training the network. The RGB image corresponds to i, g, and r bands and has been scaled with a Lupton asinh scaling (Lupton et al., 2004)
@@
The image simulation software, imSim, uses GalSim (Rowe et al., 2015) and includes physical effects due to the atmosphere and models LSST sensor features such as brighter-fatter and tree-ring effects. For a complete description, see LSST Dark Energy Science Collaboration (LSST DESC) et al. (2021). Systematic effects such as blending and variations in observing conditions are present in the simulated images and propagate into aperture photometry measurements. An example image is shown in Figure 1
@@
Using the method described above, each image contains roughly 500-700 true objects. Distributions of object redshifts are shown in Figure 2. We further limit the sample for redshift estimation to the LSST “gold" sample, defined by sources with i-band magnitude < 25.3 mag. In total, the redshift estimation branch of the network is trained with 161,205 objects. We set aside 48 images in the training set as a validation set. The test set is composed of 1925 images of 1050 pixels square. Once trained, DeepDISC can be applied to images of any size, provided enough compute resources. The details of the redshift estimation are described in the following section.
@@ 3 Methods
DeepDISC [^3] https://github.com/grantmerz/deepdisc is implemented with PyTorch (Paszke et al., 2017) and the detectron2 repository (Wu et al., 2019). We choose as baseline backbone network a Multi-scale Vision Transformer (MViTv2; Li et al., 2022) with a feature pyramid architecture (Lin et al., 2016) and cascade Region of Interest (ROI) heads (Cai & Vasconcelos, 2018). The feature pyramid network extracts features from different resolution scales in the input image, allowing for increased flexibility with objects of different sizes. Features are then sent to the Region Proposal Network (RPN) to predict where objects are in the image, and then the features corresponding to each region are pooled and sent to the ROI heads. A cascade ROI head iterates the detection stage at increasing confidence thresholds in order to produce higher quality detection inferences. A diagram of the DeepDISC architecture is shown in Figure 3
@@
We adopt the same transfer learning strategy as in Merz et al. (2023) by using a MViTv2 backbone (Li et al., 2022) with weights that have been initialized from pre-training on the ImageNet1k (Deng et al., 2009) dataset of terrestrial images. Merz et al. (2023) found that these backbones produced the best results for object detection, deblending, and star-galaxy classification, and were more robust to contrast scalings. For our initial training steps, we set the learning rate to 0.001 and train the entire network without fine-tuning, i.e., no freezing of the backbone weights. We train for 50 total epochs, lowering the learning rate by a factor of 10 each time at 15, 25, and 35 epochs. A common strategy in transfer learning is to fine-tune, or freeze the deep layers of the network and only train the first few layers on the new data. Fine-tuning is commonly used in studies where the number of dataset channels in the backbone pre-training phase matches the number of dataset channels in the full training phase. However, the 6-channel ugrizy astronomical images we use to train the network are very different from the ImageNet RGB images of everyday objects and scenes. The MViTv2 backbone must learn to extract this information to help the rest of the network. Therefore, we do not fine-tune, and instead let all layers of the network learn during training. It takes 6.5 hours to train DeepDISC on 4 NVIDIA V100 GPUs using our training set of images.
@@ 3.1 RAIL integration
In order to interface DeepDISC to the existing DESC photo-z ecosystem, we create rail_deepdisc  [^4] https://github.com/LSSTDESC/rail_deepdisc, an open-source repo that interfaces DeepDISC with the DESC software Redshift Assessment Infrastructure Layers (LSST-DESC PZ WG, 2023). RAIL is designed for end-to-end photo-z pipeline testing, with modules for creating data, applying photo-z estimators and evaluating their performance. We design rail_deepdisc to live under the RAIL ecosystem, meaning that once rail is installed, a user can import rail_deepdisc as an optional dependency along with other photo-z estimators. rail_deepdisc includes a streamlined configuration file API, interfaces with existing RAIL code, and fully parallel training and testing capabilities. Both rail_deepdisc and DeepDISC have been integrated with the LINCC Frameworks Python Project Template Oldag et al. (2024), which includes continuous integration tools for code tests, coverage, documentation, and package installation. We encourage the use of rail_deepdisc as a guide to implement other image-based codes within RAIL.
@@ 3.2.1 BPZ
BPZ is a template-based code which uses a forward model to determine the likelihood of a galaxy’s colors given an input set of template SEDs and a redshift grid. For BPZ photo-zs, We use a set of 200 template SEDs specialized for the simulated DC2 data. We note that the simulated internal galactic dust extinction contained an underlying unphysical feature in its parametrization which the principle component analysis method used to construct the SED templates was not able to capture. This feature manifests mostly at $z>1.5$, and thus high redshift performance of the BPZ algorithm will not be optimal at those redshifts, as the theoretical SEDs will not match those of the observed data. BPZ defines the posterior probability of a galaxy having redshift $z$ as $p(z|C,m_{0})$ where $C$ is the galaxy colors obtained from the observed magnitudes and m0 is the apparent magnitude in a single band (here, i-band). The posterior is obtained by marginalizing over a set of template SEDs
@@
where $p(z|T,m_{0})$ is a prior probability of a galaxy having redshift $z$ given its apparent magnitude and template type $T$. This prior is parametrized in Benítez (2000) and is determined from a best-fit to the data. This training of the prior, or “informing" the algorithm, makes BPZ a competitive template-fitting code. The BPZ model is informed with our input training catalog on a single CPU in 35 minutes.
@@ 4 Results
Figure 4: Redshift point estimates of DeepDISC (left), BPZ (middle) and FlexZBoost (right) compared to the true redshift. We use the mode of each photo-z PDF as our point estimate. The dashed line is along $z_{\rm true}=z_{\rm phot}$, and the solid lines define the $3\sigma_{{\rm IQR}}$ outlier boundary. Color corresponds to the number density of objects in each bin. Bias, scatter and outlier fraction (Equations 3, 4 and 5, respectively) are shown in the legends.
@@
In this section, we compare all codes and quantify the quality of our photo-z estimation results using several metrics designed to measure the accuracy of point estimates derived from the PDFs. After training DeepDISC, we apply the model to our test set of images. This takes about 40 minutes to run on a single NVIDIA V100 GPU and produces a catalog of 3,177,538 objects. We compute object locations using the centers of predicted bounding boxes, and use each image’s world coordinate system (WCS) to convert these pixel values to RA and DEC coordinates. In order to compare to the catalog-based codes, we cross-match to the DC2 object catalog by taking the closest gold sample match within 1". This produces a test set of 1,254,379 objects for comparison. We then obtain BPZ and FZB PDFs by applying these models to the cModel magnitudes and errors (corrected for Milky Way dust reddening) of the matched test set. As with the training set, the $1\sigma$ limiting magnitude is used for instances of non-detections or pipeline failures. We parallelize BPZ and FZB estimation over 16 CPUs. BPZ takes 31.64 minutes and FZB takes 5.08 minutes to estimate the photo-z PDFs of all test set objects.
@@
At low redshifts, DeepDISC and FZB perform the best in terms of bias and scatter. Despite this region being underrepresented in the training as there are relatively few low-z objects in the training sample N(z) (see Figure 2), DeepDISC is able to perform on par with or better than FlexZBoost. BPZ performs the worst at this low redshift range. This is likely due to a degeneracy between the Balmer break at low-z and Lyman-alpha break at high-z causing confusion between high and low-z objects, as seen in Figure 4.
@@
BPZ and FZB see a decrease in scatter beyond $\sim$2.5, at which the Lyman-alpha break enters the u-band. However, all codes underpredict the redshifts at high-z. DeepDISC is especially susceptible to this effect, and does not predict any redshifts above $\sim$2.5. During training, there is on average less than one gold sample object with z>2.5 per image. Many different attempts to address this issue were made, including weighting the redshift loss function in proportion to the redshift distribution N(z), resampling predicted regions produced by the Region Proposal Network to a more uniform redshift distribution, and doubling the training set size and number of epochs. It appears that the information that DeepDISC is able to extract from the images of sources is more meaningful for photo-z estimation when compared to source photometry. However, it is susceptible to the number of objects at a given redshift in the training set.
Point estimate metrics as a function of i-band cModel magnitude are shown in the right panel of Figure 4. In general, scatter and outlier fraction increase with magnitude for all codes. This is expected, as the signal-to-noise for faint objects is small. DeepDISC and FZB tend to perform better than BPZ at bright magnitudes despite the smaller sample size of objects. BPZ is not susceptible to this effect, but performs poorly likely due to the aforementioned Balmer and Lyman-alpha break degeneracy at low-redshift. Bright objects may also result in pipeline failures and thus the placeholder value for their magnitudes and magnitude errors causes the likelihood fitting to fail. DeepDISC is able to extract features from the images that strongly correlate with redshift in the high signal-to-noise regime.
@@
Figure 7: Recovery fraction, i.e., fraction of outliers that are no longer considered outliers if their PDF is evaluated at the secondary peak rather than the primary. The grey line at 0.515 is the result evaluated for the whole sample, and the green line is plotted vs the weight of the secondary peak. Recovery fraction varies about the mean for secondary peak weights up to about 0.4, and then drops to 0.1. The high recovery fractions indicate that the PDFs are encoding the true redshift even if the primary peak is not a reliable point estimate.
@@ 4.1 PDF metrics
An ideal PIT histogram is a uniform distribution, as true redshifts should be randomly sampled from their underlying distributions and thus give uniform CDFs. Overly narrow (overconstrained) PDFs will tend to produce a PIT histogram with an abundance of values at 0 and 1, as there is no probability mass in the region of the true redshift. Overly broad (underconstrained) PDFs tend to contain near equal amounts of probability for redshifts $z<z_{{\textrm{true}}}$ and $z>z_{{\textrm{true}}}$, and thus produce a PIT histogram with a bulge around 0.5. The ideal PIT histogram follows a uniform distribution, indicating neither too broad nor too narrow uncertainties. Systematic biases can be seen as deviations from uniformity. The quality of the PDFs can be seen in Figure 6. DeepDISC produces the least outliers at PIT values close to 0 and 1, and appears to produce the most uniform PIT histogram. As a further test of the PDF quality, we calculate the CDE loss, Kolmogorov-Smirnov (KS), Cramer von Mises (CvM), and Anderson-Darling (AD) statistics for each code. The CDE loss is estimated (up to a constant) by
@@
To further investigate the reliability of DeepDISC PDF estimates, we examine the multimodal nature of the PDFs encoded by the MDN. Since DeepDISC PDFs are parametrized by a Gaussian mixture model, we can easily examine secondary peaks. In Figure 7 we show the recovery fraction, or the fraction of outliers that no longer qualify as outliers if $z_{{\textrm{phot}}}$ is taken to be at the secondary peak value of the PDF rather than the first peak (i.e., mode). The recovery fraction for the entire test set sample is 51.15%, indicating that for a majority of outliers the secondary peak of the PDF is close to the true redshift. The secondary peak appears to reliably capture the true redshift in cases where it is both relatively strong and weak compared to the primary, as it varies about the mean until secondary peak weights above 0.4. Secondary peak weights above 0.4 indicate an extremely bimodal PDF, and the low recovery fraction indicates neither peak captures the true redshift.
In the comparisons presented above, DeepDISC photo-z outperforms traditional photo-z codes in every metric used. We have shown that among all codes, point estimates derived from DeepDISC are the most accurate and produce the least scatter and outliers, although FZB is competitive. Going beyond single point metrics, we examine the quality of the errors encoded in the PDFs by looking at the PIT histograms and additional metrics designed to measure how well-calibrated our ensemble of PDFs is. We again find that DeepDISC produces a PIT distribution closest to the ideal uniform, indicating that errors are neither too broad nor too narrow. We also find that a majority of outliers have coverage at the secondary peak of their PDFs, suggesting that our model captures physically meaningful degeneracies in its errors. In the next sections, we examine possible limiting factors of our model and dependencies of our results. We examine the effects of imaging systematics, the dependence on data quality, and possible scaling laws in the context of our model.

### v2

@@ 1 Introduction
This paper is organized as follows. In §2, we describe the simulated data and pre-processing steps. In §3, we describe our model and training methodology as well as the other photo-z estimators used in our analysis. We present the results of our trained photo-z code in §4, and quantify our photo-z estimation performance using various metrics. We discuss the results along with limitations and future applications in §5. In §6, we contextualize our findings and conclude.
@@ 2 Data
Figure 1: An example DC2 image used for training the network. The RGB image corresponds to i, r and g bands and has been scaled with a Lupton asinh scaling (Lupton et al., 2004) for visualization purposes.
@@
The image simulation software, imSim, uses GalSim (Rowe et al., 2015) and includes physical effects due to the atmosphere and models LSST sensor features such as brighter-fatter and tree-ring effects. For a complete description, see LSST Dark Energy Science Collaboration (LSST DESC) et al. (2021). Systematic effects such as blending and variations in observing conditions are present in the simulated images and propagate into aperture photometry measurements. An example image is shown in Figure 1.
@@
Using the method described above, each image contains roughly 500-700 true objects. Distributions of object redshifts are shown in Figure 2. We further limit the sample for redshift estimation to the LSST “gold” sample, defined by sources with i-band magnitude < 25.3 mag. In total, the redshift estimation branch of the network is trained with 161,205 objects. We set aside 48 images in the training set as a validation set. The test set is composed of 1925 images of 1050 pixels square. Once trained, DeepDISC can be applied to images of any size, provided enough compute resources. The details of the redshift estimation are described in the following section.
@@ 3 Methods
DeepDISC [^3] https://github.com/grantmerz/deepdisc is implemented with PyTorch (Paszke et al., 2017) and the detectron2 repository (Wu et al., 2019). We choose as baseline backbone network a Multi-scale Vision Transformer (MViTv2; Li et al., 2022) with a feature pyramid architecture (Lin et al., 2016) and cascade Region of Interest (ROI) heads (Cai & Vasconcelos, 2018). The feature pyramid network extracts features from different resolution scales in the input image, allowing for increased flexibility with objects of different sizes. Features are then sent to the Region Proposal Network (RPN) to predict where objects are in the image, and then the features corresponding to each region are pooled and sent to the ROI heads. A cascade ROI head iterates the detection stage at increasing confidence thresholds in order to produce higher quality detection inferences. A diagram of the DeepDISC architecture is shown in Figure 3.
Object detections are handled by the RPN and ROI heads. The RPN proposes a set of bounding boxes as object detections. Proposed bounding boxes are matched to the ground truth bounding boxes using an intersection-over-union (IOU) threshold. Perfectly overlapping detected/truth boxes yield an IOU of 1, those with no overlap at all yield an IOU of 0. We use the detectron2 standard IOU thresholds of [0.3,0.7], where predicted objects with a ground truth IOU less than 0.3 are considered "background", i.e., not a detection, those with 0.3<IOU<0.7 are ignored, and those with IOU>0.7 are considered to be a positive detection. Highly blended objects may result in a mismatched detection during training if the IOU criterion is met, but the goal of the RPN and ROI heads are to refine these detections during training, with the provided truth information. For more details on the RPN detection process, we refer the reader to Ren et al. (2015) and He et al. (2017).
@@
We adopt the same transfer learning strategy as in Merz et al. (2023) by using a MViTv2 backbone (Li et al., 2022) with weights that have been initialized from pre-training on the ImageNet1k (Deng et al., 2009) dataset of terrestrial images. Pre-training is often employed to speed up network convergence and reduce training costs. While the ImageNet data is very different from astronomical scenes, Burke et al. (2019) and Merz et al. (2023) show that transfer learning is a viable strategy in training instance segmentation models to infer object properties in simulated and real astronomical data. We discuss the implications and potential drawbacks of pre-training with ImageNet more in Section 4.
Burke et al. (2019) and Merz et al. (2023) used astronomical images with 3 filters, testing knowledge transfer from the RGB to another 3-color space. Photo-z estimation should use all available filters to sample object SEDs across the widest range of wavelengths possible. In this work, we use all available filters for DC2 (ugrizy) and thus test the viability of transfer learning to new color domains. In principle, DeepDISC can be applied to data with different filter sets, which in practice amounts to changing a single user-defined parameter before training.
Our choice of MViTv2 backbone network is motivated by the comparison study in Merz et al. (2023). This backbone produced the best results for object detection, deblending, and star-galaxy classification, and was more robust to contrast scalings. Contrast scaling is a commonly employed strategy to reduce the dynamic range of images input to a neural network (González et al., 2018). Astronomical images can have dynamic ranges that span several orders of magnitude, and so common strategies are to normalize images so that pixels fall in the range [0,1] and/or apply scaling functions such as an asinh scaling (Lupton et al., 2004) to ensure that faint extended regions of objects are not ignored due to very bright central regions. Merz et al. (2023) found the MViTv2 backbone to be very robust to different contrast scalings for instance segmentation tasks. Networks trained with scalings with large or small dynamic ranges showed little difference in performance. Based on those results, we apply no scaling to the DC2 data and use the raw images. This poses the advantage of maintaining noise and background properties that may otherwise be altered through changing the dynamic range.
For our initial training steps, we set the learning rate to 0.001 and train the entire network without fine-tuning, i.e., no freezing of the backbone weights. We train for 50 total epochs, lowering the learning rate by a factor of 10 each time at 15, 25, and 35 epochs. A common strategy in transfer learning is to fine-tune, or freeze the deep layers of the network and only train the first few layers on the new data. Fine-tuning is commonly used in studies where the number of channels (colors) in the dataset used in the pre-training phase matches the number of dataset channels in the full training phase. However, the 6-channel ugrizy astronomical images we use to train the network are very different from the ImageNet RGB images of everyday objects and scenes. The MViTv2 backbone must learn to extract this information to help the rest of the network. Therefore, we do not fine-tune, and instead let all layers of the network learn during training. It takes 6.5 hours to train DeepDISC on 4 NVIDIA V100 GPUs using our training set of images.
@@ 3.1 RAIL integration
In order to interface DeepDISC to the existing DESC photo-z ecosystem, we present rail_deepdisc [^4] https://github.com/LSSTDESC/rail_deepdisc, an open-source repo that interfaces DeepDISC with the DESC software Redshift Assessment Infrastructure Layers (LSST-DESC PZ WG, 2023). RAIL is designed for end-to-end photo-z pipeline testing, with modules for creating data, applying photo-z estimators and evaluating their performance. We design rail_deepdisc to live under the RAIL ecosystem, meaning that once rail is installed, a user can import rail_deepdisc as an optional dependency along with other photo-z estimators. rail_deepdisc includes a streamlined configuration file API, interfaces with existing RAIL code, and fully parallel training and testing capabilities. Both rail_deepdisc and DeepDISC have been integrated with the LINCC Frameworks Python Project Template Oldag et al. (2024), which includes continuous integration tools for code tests, coverage, documentation, and package installation. We encourage the use of rail_deepdisc as a guide to implement other image-based codes within RAIL.
@@ 3.2 Code Comparison
Both catalog-level codes are informed by the same training data set. In order to conduct a fair comparison with DeepDISC, we take the set of 161,205 gold sample entries in the truth catalog used by DeepDISC and take the corresponding matched set of entries in the object catalog. The object catalog is produced by running the LSST Science pipelines software on the simulated images, which detects and deblends objects (LSST Dark Energy Science Collaboration (LSST DESC) et al., 2021). We train BPZ and FZB using the object catalog as this incorporates blending and other imaging systematics, which DeepDISC encounters at the image level. This ensures all methods are informed with as close to the same prior information as possible. However, we note that due to a flaw in the dust extinction law assumed in the simulations, the colors of our simulated galaxies are not fully represented by the template set employed by BPZ. This mismatch, along with the presence of imaging systematics in the derived photometric catalog will very likely degrade the performance of BPZ relative to Schmidt et al. (2020).
In order to assign redshifts to the object catalog to use for training, we use the DC2 truth_match catalogs. Entries in the truth catalog are matched to entries in the object catalog by first taking all truth entries within 1′′ of an object entry, and within an r-band magnitude difference of 1 mag. Then, the truth entry with the smallest r-band magnitude difference is matched to a object entry. If no truth entry meets this criterion, the closest truth entry is matched. We also filter by selecting truth_match entries with the flag detect_isPrimary=True. This flag helps to produce a catalog of unique objects, as it filters out duplicate objects appearing in the overlapping region of multiple patches. It also filters out isolated objects that had the deblender applied to them, in favor of using nondeblended measurements for isolated objects.
As with DeepDISC photo-zs, we limit the training of BPZ and FZB to objects with an i-band magnitude of less than 25.3 mag. We use forced photometry cModel magnitudes and magnitude errors as inputs for both BPZ and FZB, and account for Milky Way dust reddening by applying a magnitude correction in each band
$mag_{\rm dereddened}=mag-R_{\lambda}*E(B-V)$ (1)
[…]
@@ 3.2.1 BPZ
BPZ is a template-based code which uses a model to determine the likelihood of a galaxy’s colors given an input set of template SEDs and a redshift grid. For BPZ photo-zs, We use a set of 200 template SEDs specialized for the simulated DC2 data. We note that the simulated internal galactic dust extinction contained an underlying unphysical feature in its parametrization which the principle component analysis method used to construct the SED templates was not able to capture. This feature manifests mostly at $z>1.5$, and thus high redshift performance of the BPZ algorithm will not be optimal at those redshifts, as the theoretical SEDs will not match those of the observed data.
The following notation uses magnitudes, as is the original convention from Benítez (2000). However, we note that the BPZ algorithm converts input magnitudes to fluxes that are more well-behaved and have Gaussian errors. BPZ defines the posterior probability of a galaxy having redshift $z$ as $p(z|C,m_{0})$ where $C$ is the galaxy colors obtained from the observed magnitudes and m0 is the apparent magnitude in a single band (here, i-band). The posterior is obtained by marginalizing over a set of template SEDs
@@
where $p(z|T,m_{0})$ is a prior probability of a galaxy having redshift $z$ given its apparent magnitude and template type $T$ and $p(C|z,T)$ is the likelihood of a galaxy having colors $C$ given a redshift and template type. This prior is parametrized in Benítez (2000) and is determined from a best-fit to the data. This training of the prior, or “informing" the algorithm, makes BPZ a competitive template-fitting code. For more detailed functional forms of the likelihood and prior, see Benítez (2000). The BPZ model is informed with our input training catalog on a single CPU in 35 minutes.
@@ 4 Results
Figure 4: Redshift point estimates of DeepDISC (left), BPZ (middle) and FlexZBoost (right) compared to the true redshift. We use the mode of each photo-z PDF as our point estimate. The dashed line is along $z_{\rm true}=z_{\rm phot}$, and the solid lines define the $3\sigma_{\rm IQR}$ outlier boundary. Color corresponds to the number density of objects in each bin. Bias, scatter and outlier fraction (Equations 4, 5 and 6, respectively) are shown in the legends.
@@
In this section, we compare all codes and quantify the quality of our photo-z estimation results using several metrics designed to measure the accuracy of point estimates derived from the PDFs. After training DeepDISC, we apply the model to our test set of images. This takes about 40 minutes to run on a single NVIDIA V100 GPU and produces a catalog of 3,177,538 objects. We compute object locations using the centers of predicted bounding boxes, and use each image’s world coordinate system (WCS) to convert these pixel values to RA and DEC coordinates. In order to compare to the catalog-based codes, we cross-match to the DC2 object catalog by taking the closest gold sample match within 1′′. This produces a test set of 1,254,379 objects for comparison. This detection catalog is 94.2% complete with gold sample objects in the DC2 truth catalog and 96.9% complete with the gold sample of the DC2 object catalog. We find that DeepDISC detection completeness tends to decrease with brighter i-band magnitudes, which may be due to both overdeblending, or large, bright objects being truncated by the image cutout.
We then obtain BPZ and FZB PDFs by applying these models to the cModel magnitudes (corrected for Milky Way dust reddening) and errors of the matched test set. As with the training set, the $1\sigma$ limiting magnitude is used for instances of non-detections or pipeline failures. We parallelize BPZ and FZB estimation over 16 CPUs. BPZ takes 31.64 minutes and FZB takes 5.08 minutes to estimate the photo-z PDFs of all test set objects.
@@
At low redshifts, DeepDISC and FZB perform the best in terms of bias and scatter. Despite this region being underrepresented in the training as there are relatively few low-z objects in the training sample N(z) (see Figure 2), DeepDISC is able to perform on par with or better than FlexZBoost. BPZ performs the worst at this low redshift range. This is likely due to a degeneracy between the Balmer break at low-z and Lyman break at high-z causing confusion between high and low-z objects, as seen in Figure 4.
@@
BPZ and FZB see a decrease in scatter beyond $\sim$2.5, at which the Lyman break enters the u-band. However, all codes underpredict the redshifts at high-z. Some poor performance at the edges of the redshift distribution is expected, due to noise and the choice of redshift grid used to parametrize the PDFs. Additionally, mode point estimates fail in the case of highly bimodal PDFs (high redshift blue galaxies may be confused for low redshift red galaxies). However, DeepDISC is especially sensitive to this regime, as no photo-z PDFs produced by DeepDISC yield a mode above $\sim$2.5. During training, there is on average less than one gold sample object with z>2.5 per image. Many different attempts to address this issue were made, including weighting the redshift loss function in proportion to the redshift distribution N(z), resampling predicted regions produced by the Region Proposal Network to a more uniform redshift distribution, and doubling the training set size and number of epochs. It appears that the information that DeepDISC is able to extract from the images of sources is more meaningful for photo-z estimation when compared to source photometry. However, it is susceptible to the number of objects at a given redshift in the training set. We discuss this result and possible solutions more in Section 5.
Point estimate metrics as a function of i-band cModel magnitude are shown in the right panel of Figure 4. In general, scatter and outlier fraction increase with magnitude for all codes. This is expected, as the signal-to-noise for faint objects is small. DeepDISC and FZB tend to perform better than BPZ at bright magnitudes despite the smaller sample size of objects. The larger bias of BPZ photo-zs at bright magnitudes reflect a phenomenon with template fitting codes due to the smaller photometric uncertainties measured at bright magnitudes. Smaller errors can lead to larger $\chi^{2}$ values in the likelihood fitting process which exaggerate the importance of biased flux values. This will be especially noticeable if the photometric errors measured by the LSST science pipeline are underestimated.
@@
Figure 7: Recovery fraction, i.e., fraction of outliers that are no longer considered outliers if their PDF is evaluated at the secondary peak rather than the primary. We only include secondary peaks with a weight of at least 0.1 in the Figure, and plot the peak weight on the x-axis. The high recovery fractions indicate that the PDFs are encoding the true redshift even if the primary peak is not a reliable point estimate.
@@ 4.1 PDF metrics
An ideal PIT histogram is a uniform distribution, as true redshifts should be randomly sampled from their underlying distributions and thus give uniform CDFs. Overly narrow (overconstrained) PDFs will tend to produce a PIT histogram with an abundance of values at 0 and 1, as there is no probability mass in the region of the true redshift. Overly broad (underconstrained) PDFs tend to contain near equal amounts of probability for redshifts $z<z_{\textrm{true}}$ and $z>z_{\textrm{true}}$, and thus produce a PIT histogram with a bulge around 0.5. The ideal PIT histogram follows a uniform distribution, indicating neither too broad nor too narrow uncertainties. Systematic biases can be seen as deviations from uniformity. The quality of the PDFs can be seen in Figure 6. DeepDISC produces the least outliers at PIT values close to 0 and 1, and appears to produce the most uniform PIT histogram. The BPZ PIT histogram is noticeably overconstrained, indicating BPZ tends to underestimate the uncertainty on its redshift estimates. This may also be due to underestimated photometric uncertainties which are measured from the simulated coadded images using the LSST Science pipelines software.
As a further test of the PDF quality, we calculate the CDE loss, Kolmogorov-Smirnov (KS), Cramer von Mises (CvM), and Anderson-Darling (AD) statistics for each code. The CDE loss is estimated (up to a constant) by
@@
To further investigate the reliability of DeepDISC PDF estimates, we examine the multimodal nature of the PDFs encoded by the MDN. We use a peak finding algorithm to determine secondary modes of the PDFs, and examine whether the secondary modes capture meaningful information. We find that 23.5% of the DeepDISC PDFs have secondary peaks. We define the relative peak weight for each PDF $p(z)$ by $\frac{p(z_{2})}{\sum_{i}p(z_{i})}$ where $z_{i}$ are the peaks. 13.3% of DeepDISC PDFs have a secondary peak with a weight above 0.1. For BPZ, 61.5% of the PDFs have a secondary peak and 37.9% of the PDFs have a weight above 0.1. FZB yields 52.4% of PDFs with secondary peaks and 43.7% with the peak weight above 0.1. Given the relatively simple nature of the DeepDISC Gaussian mixture model parametrization, a lower number of secondary peaks compared to the other codes is not unexpected. However, to see whether these secondary modes contain meaningful information, we evaluate the recovery fraction for these PDFs. The recovery fraction is defined as the fraction of photo-z outliers that no longer qualify as outliers if $z_{\textrm{phot}}$ is taken to be at the secondary mode of the PDF rather than the first mode. We show the recovery fraction as a function of secondary peak weights in Figure 7, limiting the analysis to PDFs with secondary peaks with a weight above 0.1. For all codes, recovery fractions are above 50% for strong and weak peaks, indicating that the photo-z estimators are able to capture meaningful degeneracies in the PDFs.
Overall, we have shown that under conditions of closely matched prior information, is very competitive in its photo-z estimation. In our experimental setup, DeepDISC photo-z outperforms traditional photo-z codes in terms of bias, scatter, and outlier fraction. Going beyond single point metrics, we examine the quality of the errors encoded in the PDFs by looking at the PIT histograms and additional metrics designed to measure how well-calibrated our ensemble of PDFs is. We again find that DeepDISC produces a PIT distribution closest to the ideal uniform, indicating that errors are neither too broad nor too narrow. We also find that for all codes, a majority of outliers have coverage at the secondary peak of their PDFs, suggesting that the models captures physically meaningful degeneracies. In the next Sections, we examine possible limiting factors of our model and dependencies of our results. We examine the effects of imaging systematics, the dependence on data quality, and possible scaling laws in the context of our model.